OpenStudy (anonymous):

for questions 5 and 6 find the slope of the line that goes through the following points. Will give medal

OpenStudy (4everaddicted2anime):

What are the points for 5 and 6?

OpenStudy (anonymous):

Asap please 😊

OpenStudy (anonymous):

I dont know, that picture is all it gives

OpenStudy (4everaddicted2anime):

There is no picture

OpenStudy (anonymous):

Here

OpenStudy (4everaddicted2anime):

Okay so for number 5 lets use the points (0,-4) and (1,-3). The slope formula is \[\frac{ y_2-y_1 }{ x_2-x_1 }\] Plug in the values for x and y and you should get\[\frac{ -3--4 }{ 1-0 }\] After simplifying you should get \[\frac{ 1 }{ 1 }\] This simply equals 1. The answer to number 5 is 1.

OpenStudy (4everaddicted2anime):

For number 6 lets use the points (-1,-4) and (0,-1). using the same slope formula as above plug in the values for x and y respectively to get \[\frac{ -1--4 }{ 0--1 }\] After simplifying you should get \[\frac{ 3 }{ 1 }\] This simply equals 3. The answer to number 6 is 3.

OpenStudy (anonymous):

Thank you so much !!!

OpenStudy (4everaddicted2anime):

yw